OpenStudy (anonymous):

How do you do -13=5(1+4m)- 2m. I already simplified and got -13=5+18m

OpenStudy (anonymous):

first distribute -13=5+20m-2m now combine like terms -13= 5+18m now subtract 5 from both sides -18=18m solution: m=-1

OpenStudy (anonymous):

ok. Thank you.

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

do i get a medal?

OpenStudy (anonymous):

You should get one